# Beyond the Use-and-then-Forget (UatF) Bound: Fixed Point Algorithms for Statistical Max-Min Power Control

### v1

@@ Abstract
We introduce mathematical tools and fixed point algorithms for optimal statistical max-min power control in cellular and cell-less massive MIMO systems. Unlike previous studies that rely on the use-and-then-forget (UatF) lower bound on Shannon achievable (ergodic) rates, our proposed framework can deal with alternative bounds that explicitly consider perfect or imperfect channel state information (CSI) at the decoder. In doing so, we address limitations of UatF-based algorithms, which inherit the shortcomings of the UatF bound. For example, the UatF bound can be overly conservative: in extreme cases, under fully statistical (nonadaptive) beamforming in zero-mean channels, the UatF bound produces trivial (zero) rate bounds. It also lacks scale invariance: merely scaling the beamformers can change the bound drastically, especially when simple beamforming strategies are employed. In contrast, our framework is compatible with information-theoretic bounds that do not suffer from the above drawbacks. We illustrate the framework by solving a max-min power control problem considering a standard bound that exploits instantaneous CSI at the decoder.
@@ I Introduction
The use-and-then-forget (UatF) bound has been fundamental in the development of beamforming and power control algorithms in massive MIMO and cell-less networks [1, 2, 3, 4, 5, 6, 7] because it often leads to tractable and scalable optimization problems. Furthermore, in some scenarios of practical relevance – e.g., when optimal beamformers are employed – the UatF bound often provides spectral efficiency predictions that closely match those obtained from tighter bounds of the Shannon ergodic rates, including the coherent decoding bound [1, 7].
@@
Against this background, we demonstrate that, in networks where the beamforming strategy is independent of the transmit power vector (the general case will be considered elsewhere), alternatives to the UatF bound – e.g., the coherent decoding bound and the upper bound in [8, Proposition 1] – can be employed in tractable optimization problems, provided we accept that expectations are approximated via Monte Carlo sampling, as also done in power control problems based on the UatF bound [2, 8, 7, 9]. Nevertheless, we avoid additional approximations, such as replacing the expectation of a ratio with the ratio of expectations.
@@ II-B Framework for Power Control with Tight Ergodic Rate Bounds
Although the above aspects might prevent the direct application of some results for standard interference functions to MSP functions – because, as discussed above, the set of continuous standard interference functions restricted to the positive cone is a proper subset of MSP functions [^2] In turn, MSP functions are a proper subclass of order-preserving subhomogeneous functions [18, 20]. –, many key properties remain analogous. In particular, the class of MSP functions is closed under addition, positive scaling, and finite maxima and minima, just as in the case of standard interference functions. Furthermore, any MSP mapping $T:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}$ is contractive in the complete metric space $({\mathbb{R}}_{++}^{N},d_{\mathrm{T}})$ [18, Lemma 2.1.7] (but not necessarily a Lipschitz contraction), where $d_{\mathrm{T}}:{\mathbb{R}}_{++}^{N}\times{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}$ is Thompson’s metric defined in [18, p. 30], so $T$ has at most one fixed point. If $\mathrm{Fix}(T)\neq\emptyset$, arguments similar to those described in [10] to prove convergence of the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via the fixed point iterations $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=T({\bm{x}}_{n})$, with ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, can be used for MSP mappings. In addition, the following property, which is crucial for the power control algorithms derived in the next section, can be established using the same arguments in [11], and they are therefore omitted.
@@ III Power control in multi-user MIMO
The beamformers in (5) are random vectors, so they are allowed to change for each channel realization. However, as clear from the notation, they should not depend on the power vector $\bm{p}$}, the argument of the functions $(r_{u})_{u\in\mathcal{U}}$. Prominent beamforming strategies satisfying this assumption include conventional conjugate beamforming, (regularized) zero-forcing beamforming, and MMSE-like beamformers such as those simulated in [9] and the references therein. The general case will be considered elsewhere.
@@
We can use the same arguments in [5, Proposition 2] with standard interference functions replaced with MSP functions to show that, although the solution to (16) (and, hence, (12)) may not be unique, there always exists a solution achieving equality in all constraints; i.e., there exists a solution $(\gamma^{\star},{\bm{p}}^{\star})$ satisfying $(\gamma^{\star})^{-1}{\bm{p}}^{\star}=T({\bm{p}}^{\star})$ and $(1/p_{\text{max}})\|{\bm{p}}^{\star}\|=1$. These equalities establish that $(\gamma^{\star})^{-1}$ and $\bm{p}$$^{\star}$ are, respectively, the conditional eigenvalue and conditional eigenvector of the mapping $T$ and the monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|_{\star}:=(1/p_{\text{max}})\|{\bm{p}}\|$. Therefore, as a corollary of Proposition 1, Proposition 3, and [5, Proposition 2] (after adapting the proof of this last proposition by substituting MSP functions with standard interference functions), we obtain the following result, which shows a simple fixed point method that provably converges to the solution to (12) with minimum total power:
@@
Let the assumptions in Proposition 3 be valid. Denote by $\mathcal{S}\neq\emptyset$ the set of solutions to (12) and let $\|\cdot\|$ be the monotone norm in (12). Given ${\bm{p}}_{1}\in{\mathbb{R}}_{++}^{N}$, the sequence generated $({\bm{p}}_{n})_{n\in{\mathbb{N}}}$ via $(\forall n\in{\mathbb{N}})~{\bm{p}}_{n+1}=(p_{\mathrm{max}}/\|T({\bm{p}}_{n})\|)T({\bm{p}}_{n})$ converges to the power vector ${\bm{p}}^{\star}\in{\mathbb{R}}_{++}^{N}$ that solves the optimization problem
@@ IV Simulations
The performance evaluation considers uniform weights in (12) (i.e., $(\forall u\in\mathcal{U})~\alpha_{u}=1$) and centralized cell-less MMSE beamforming design with fixed (unitary) power weighting coefficients in the channel inversion stage [9, Sect. 3-D]. We estimate the ergodic rates by averaging across 500 Monte Carlo trials, each drawing an independent channel realization for a single, fixed network layout. In the proposed fixed point iterations, shown in Corollary 1, we use (5) as the utility function, which is called “optimistic ergodic rate” (OER) in the following (see Remark 3). As a baseline, we compute the OER bound with the solution to the standard max-min problem under the UatF bound, which admits a pseudo-closed-form expression [9, Prop. 4]. This approach mirrors established practice in the literature: if simple fixed point algorithms capable of solving (12) were unknown (as was the case prior to this study, to the best of our knowledge), we would consider the surrogate max-min power-control problem that uses the UatF bound instead of the target bound in (5), with the expectation that the surrogate solution would approximate the solution to (12).

### v2

@@ Abstract
We introduce mathematical tools and fixed point algorithms for optimal statistical max-min power control in cellular and cell-less massive MIMO systems. Unlike previous studies that rely on the use-and-then-forget (UatF) lower bound on Shannon achievable (ergodic) rates, our proposed framework can deal with alternative bounds that explicitly consider perfect or imperfect channel state information (CSI) at the decoder. In doing so, we address limitations of UatF-based power control algorithms, which inherit the shortcomings of the UatF bound. For example, the UatF bound can be overly conservative: in extreme cases, under fully statistical (nonadaptive) beamforming in zero-mean channels, the UatF bound produces trivial (zero) rate bounds. It also lacks scale invariance: merely scaling the beamformers can change the bound drastically. In contrast, our framework is compatible with information-theoretic bounds that do not suffer from the above drawbacks. We illustrate the framework by solving a max-min power control problem considering a standard bound that exploits instantaneous CSI at the decoder.
@@ I Introduction
The use-and-then-forget (UatF) bound has been fundamental in the development of beamforming and power control algorithms in massive MIMO and cell-less networks because it often leads to tractable and scalable optimization problems [1, 2, 3, 4, 5, 6, 7]. Furthermore, in some scenarios of practical relevance – e.g., when optimal beamformers are employed – the UatF bound provides spectral efficiency predictions that closely match those obtained from tighter bounds of the Shannon ergodic rates, including the coherent decoding bound [1, 7].
@@
Against this background, we demonstrate that, in networks where the beamforming strategy is independent of the transmit power vector (the general case will be considered elsewhere), alternatives to the UatF bound – e.g., the coherent decoding bound and the upper bound in [8, Proposition 1] – can be employed in tractable weighted max-min optimization problems, provided we accept that expectations are approximated via Monte Carlo sampling, as also done in power control problems based on the UatF bound [2, 8, 7, 9]. Nevertheless, we avoid additional approximations, such as replacing the expectation of a ratio with the ratio of expectations.
@@ II-B Framework for Power Control with Tight Ergodic Rate Bounds
Although the above aspects might prevent the direct application of some results for standard interference functions to MSP functions – because, as discussed above, the set of continuous standard interference functions restricted to the positive cone is a proper subset of MSP functions [^2] In turn, MSP functions are a proper subclass of order-preserving subhomogeneous functions [18, 20]. –, many key properties remain analogous. In particular, the class of MSP functions is closed under addition, positive scaling, and finite maxima and minima, just as in the case of standard interference functions. Furthermore, any MSP mapping $T:{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{++}^{N}$ is contractive in the complete metric space $({\mathbb{R}}_{++}^{N},d)$ [18, Lemma 2.1.7] (but not necessarily a Lipschitz contraction), where $d:{\mathbb{R}}_{++}^{N}\times{\mathbb{R}}_{++}^{N}\to{\mathbb{R}}_{+}$ is Thompson’s metric defined in [18, p. 30], so $T$ has at most one fixed point. If $\mathrm{Fix}(T)\neq\emptyset$, arguments similar to those described in [10] to prove convergence of the sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ generated via the fixed point iterations $(\forall n\in{\mathbb{N}})~{\bm{x}}_{n+1}=T({\bm{x}}_{n})$, with ${\bm{x}}_{1}\in{\mathbb{R}}_{++}^{N}$, can be used for MSP mappings. In addition, the following property, which is crucial for the power control algorithms derived in the next section, can be established using the same arguments in [11], and they are therefore omitted.
@@ III Power control in multi-user MIMO
The beamformers $({\bm{v}}_{u}:\Omega\to\mathbb{C}^{K})_{u\in\mathcal{U}}$ in (8) are random vectors, so they are allowed to change for each channel realization. However, as clear from the notation, they should not depend on the power vector $\bm{p}$}, the argument of the functions $(r_{u})_{u\in\mathcal{U}}$. Prominent beamforming strategies satisfying this assumption include conventional conjugate beamforming, (regularized) zero-forcing beamforming, and MMSE-like beamformers such as those simulated in [9] and the references therein. The general case will be considered elsewhere.
@@
We can use the same arguments in [5, Proposition 2] with standard interference functions replaced with MSP functions to show that, although the solution to (16) (and, hence, (12)) may not be unique, there always exists a solution achieving equality in all constraints; i.e., there exists a solution $(\gamma^{\star},{\bm{p}}^{\star})$ satisfying $(\gamma^{\star})^{-1}{\bm{p}}^{\star}=T({\bm{p}}^{\star})$ and $(1/p_{\text{max}})\|{\bm{p}}^{\star}\|=1$. These equalities establish that $(\gamma^{\star})^{-1}$ and $\bm{p}$$^{\star}$ are, respectively, the conditional eigenvalue and conditional eigenvector of the mapping $T$ and the monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|_{\star}:=(1/p_{\text{max}})\|{\bm{p}}\|$. Therefore, as a corollary of Proposition 1, Proposition 3, and [5, Proposition 2] (after adapting the proof of this last proposition by substituting standard interference functions with MSP functions), we obtain the following result, which shows a simple fixed point method that provably converges to the solution to (12) with minimum total power:
@@
Let the assumptions in Proposition 3 be valid. Denote by $\mathcal{S}\neq\emptyset$ the set of solutions to (12) and let $\|\cdot\|$ be the monotone norm in (12). Given ${\bm{p}}_{1}\in{\mathbb{R}}_{++}^{N}$, the sequence $({\bm{p}}_{n})_{n\in{\mathbb{N}}}$ generated via $(\forall n\in{\mathbb{N}})~{\bm{p}}_{n+1}=(p_{\mathrm{max}}/\|T({\bm{p}}_{n})\|)T({\bm{p}}_{n})$ converges to the power vector ${\bm{p}}^{\star}\in{\mathbb{R}}_{++}^{N}$ that solves the optimization problem
@@ IV Simulations
The performance evaluation considers uniform weights in (12) (i.e., $(\forall u\in\mathcal{U})~\alpha_{u}=1$) and centralized cell-less MMSE beamforming design with fixed (unitary) power weighting coefficients in the channel inversion stage [9, Sect. 3-D]. We estimate the ergodic rates by averaging across 500 Monte Carlo trials, each drawing an independent channel realization for a single, fixed network layout. In the proposed fixed point iterations, shown in Corollary 1, we use (5) as the utility function, which is called “optimistic ergodic rate” (OER) in the following (see Remark 3). As a baseline, we compute the OER bound with the solution to the standard max-min problem under the UatF bound, which admits a pseudo-closed-form expression [9, Proposition 4]. This approach mirrors established practice in the literature: if simple fixed point algorithms capable of solving (12) were unknown (as was the case prior to this study, to the best of our knowledge), we would consider the surrogate max-min power-control problem that uses the UatF bound instead of the target bound in (5), with the expectation that the surrogate solution would approximate the solution to (12).